Given an example of an elliptic partial differential equation of 2nd order,





justifying your answer

1
Expert's answer
2022-04-14T08:17:59-0400

The Laplace's equation is the example of an elliptic partial differential equation of 2nd order

"\nabla^2\phi(x,y,z)=\frac{\partial^2\phi}{\partial x^2}+\frac{\partial^2\phi}{\partial y^2}+\frac{\partial^2\phi}{\partial z^2}=0"